OpenStudy (anonymous):

Predict the distance at which two cars will collide when they start at opposite ends separated by a distance of 4 meters. One car leaves 1.5 seconds after the first car is released. One car has a velocity of 0.38 m/s and the other has a velocity of 0.31 m/s.

OpenStudy (anonymous):

Try = 2aS=Vf^2-Vi^2

OpenStudy (anonymous):

Vf=vi+at for a

OpenStudy (anonymous):

what am i plugging in?

OpenStudy (anonymous):

Take vi=0 and put the value now you have got vf for both cars..put the velocity for the car which started 1.5 sec before and find the distance..

OpenStudy (anonymous):

You can get acceleration from vf=vi+at..then find S..where S is distance...